OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x + 2y = 6. Answer x-intercept: (0, 2), y-intercept: (0, 3) x-intercept: (2, 0), y-intercept: (0, 3) x-intercept: (3, 0), y-intercept: (2, 0) x-intercept: (0, 3), y-intercept: (0, 2)

OpenStudy (dream95):

the answer is the second one, try to substitute the value into the equation..to get the answer

OpenStudy (anonymous):

for x intercept 3x+0=6, x=2 (2,0) for y intercept 0+2y=6, y=3 (0,3)

OpenStudy (anonymous):

thanks .